OpenStudy (anonymous):

Find the equation of the plane through the point P = (2, 3, 2) and parallel to the plane −(x+4y+z)=2.

OpenStudy (amistre64):

parallel planes have parallel normals right?

OpenStudy (anonymous):

correct

OpenStudy (amistre64):

steal the normal from the other one and use it on the point given

OpenStudy (anonymous):

so <-1, 4, 1> for the normal vector?

OpenStudy (anonymous):

oops...<-1,-4,-1>

OpenStudy (amistre64):

<1,4,1> looks to suffice

OpenStudy (anonymous):

even with the negative out front?

OpenStudy (amistre64):

yep, the negative is a scalar

OpenStudy (amistre64):

it could just as easily be 6 out front

OpenStudy (anonymous):

okay...so then it goes to (x-2) +4(y-3) + (z-2) = 0?

OpenStudy (amistre64):

looks good to me, simplify as form dictates

OpenStudy (anonymous):

x - 2 +4y - 12 + z - 2 x+4y + z = -16

OpenStudy (anonymous):

x + y + z = -4?

OpenStudy (amistre64):

(x-2) +4(y-3) + (z-2) = 0 x + 4y +z -2-12-2 = 0 x + 4y +z -16 = 0 , is good

OpenStudy (anonymous):

i have to give it as z = again

OpenStudy (amistre64):

then put it all to the right but z :)

OpenStudy (anonymous):

z = -x -4y +16?

OpenStudy (amistre64):

z = 16 -4y -x

OpenStudy (anonymous):

excellent...thanx again

OpenStudy (amistre64):

youre welcome